OpenStudy (isuckatschool5236856):

Find the slope of the line whose equation is x - 6y + 12 = 0.

OpenStudy (isuckatschool5236856):

so i am IN A THING CALLED TRIPLE A THIS IS MY LAST COURSE I AM TAKING AND I REALLY SUCK AT MATH so please help me anyway possible

OpenStudy (nitroster):

x - 6y = -12 -6y = -12 - x y = 0.16x + 2

OpenStudy (isuckatschool5236856):

i want to ask y'all a question if i may? were are y'all from, like what are y'all doing are y'all teachers helping people out or or you like another student because i am a senior in high school and i just don't understand the situation on this website lol.

OpenStudy (nitroster):

Well... I just like to help people out. Even tho, my questions aren't always answered sadly :(